OpenStudy (anonymous):

Given the function f(x) = 3x, find the value of f−1(81). f−1(81) = 1 f−1(81) = 2 f−1(81) = 3 f−1(81) = 4

OpenStudy (astrophysics):

First find the inverse of 3x, know how?

OpenStudy (astrophysics):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

@D3xt3R

OpenStudy (anonymous):

@bribri50 is \[f(x)=3^x~~or~~3x\]??

OpenStudy (anonymous):

you are being asked to solve \[3^x=81\] for \(x\) it should be more or less obvious

OpenStudy (zzr0ck3r):

@D3xt3R look at the answers:)

OpenStudy (anonymous):

\[f(x)=3^x\] \[x=3^{f(x)}\] \[f^{-1}(x)=log_3(x)\] \[f^{-1}(81)=log_3(81)\] \[f^{-1}(81)=log_3(3^4)\] \[\boxed{f^{-1}(81)=4}\]

OpenStudy (astrophysics):

Oh given the options, that makes much more sense!

OpenStudy (anonymous):

;D